# A Generative Framework for Probabilistic, Spatiotemporally Coherent Downscaling of Climate Simulation

### v3

@@ Methods
The code (in Python) for model training and evaluation, data processing, and for reproducing all presented experiments and figures contained in this manuscript, is available at https://github.com/schmidtjonathan/Climate2Weather.
Appendix A Supplementary Text and Figures
A.1 Spatial region: coarse and fine grid
Figure A.1: This plot shows the spatial region considered in this study. The coarse $8\times 8$-node grid is marked with purple crosses. The fine $128\times 128$-node grid is marked with green dots. Each coarse-grid node lies in the center of a corresponding $16\times 16$-patch of high-resolution grid nodes. Both grids span exactly the same area.
A.2 Spatial patterns on multiple scales
To assess the performance of the downscaling model across multiple spatial length scales, we show that the radially-averaged power spectral densities (RAPSD) [43] of the predictions align with the reanalysis data in figure A.2. The RAPSD is computed by averaging the power spectrum over all directions of the same wavenumber in Fourier space. The quantity is commonly used in the context of weather dynamics, especially when estimating precipitation [44, 45, for example]. We use the open-source pysteps package by Pulkkinen et al. [46] to compute the RAPSD. The reported RAPSD values are averages over the considered time period.
@@ A.2 Spatial patterns on multiple scales
Figure A.2: This plot shows the RAPSD for reanalysis data, coarse input, and the corresponding downscaled predictions. Subplot (a) mirrors the experimental setup of the on-model experiment that predicts high-resolution reanalysis data during the cyclone "Friederike" in January 2018. Subplot (b) covers the CMIP6-downscaling setup as described in the Methods section.
@@
A.3 Embedding the predicted region in an extended spatial context
@@
We embed the high-resolution predictions during the cyclone "Friederike" (c.f. figure 4) into a larger spatial context in order to learn about how long-distance interconnections between the studied spatial region and its surroundings are captured by the model. We argue that it is likely that our statistical downscaling model is able to capture the global connectedness of weather dynamics that are contained in the reanalysis data it is trained on. Concretely, figures A.3 to A.6 visualize the spatiotemporal dynamics of downscaled samples, reanalysis data, and conditioning input. Thereby, the spatial region is extended beyond the one considered in this study. Reanalysis data is used to fill in the regions outside of the predicted patch. Clearly visible or implausible transitions at the edges of the patch would indicate that the model predictions do not align with the surrounding spatial context. We find, however, that there is a smooth transition from outside the predicted area to its interior for the generated high-resolution predictions. The generated local dynamics seamlessly fit into the more surrounding context of reanalysis data.
@@
Figure A.3: Predicting mean sea-level pressure during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Figure A.4: Predicting surface air temperature during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Figure A.5: Predicting zonal wind speed during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Figure A.6: Predicting meridional wind speed during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
A.4 Wind power prediction
@@
A crucial motivation for spatiotemporal downscaling climate simulations to the weather scale are downstream tasks that require future local weather patterns. We provide an exemplary evaluation of the estimated generated wind powers as computed from the CMIP6 simulation, the downscaled predictions, and the reanalysis data. We find that, when computing the spatial average of generated windpowers, the wind-speed and wind-power predictions of the ESM are matched by the downscaled samples in distribution (figure A.7). Comparing single locations reveals that the ESM locally sometimes over-predicts or under-predicts the wind power generated following the reanalysis data. For both cases, we consistently find that our model corrects the respective over- and under-estimation for multiple randomly-selected locations. We compare estimated densities of the wind-speed values (figure A.7 a) and use this distribution to derive the amount of wind power generated from the respectively predicted wind speeds (figure A.7 b). For that, we first compute the wind-power curve for a range of wind-speed values (00m/s to $30$m/s), using the open-source package windpowerlib [47] and an arbitrary wind-turbine model (turbine type: "E-115/3000", hub-height: $100$m). Then, we weigh this power curve with the density of predicted wind speeds to obtain the predicted wind power for each wind speed, taking into account how likely this wind speed value is to occur in the predictions. Finally, we compare the accumulated generated wind power over time. We show the comparison between two pairs of locations in figure A.8.
@@
Figure A.7: Estimated generated wind power over time, averaged over the spatial domain. This plot analyzes the local wind-speed predictions via the generated wind-power derived from the wind speeds. The full spatial region is considered over the year 2014. Subplot a visualizes a kernel-density estimate of the aggregated wind speeds for reanalysis data (green), ESM simulations (purple) and downscaled predictions (black). In b, we plot the generated wind power for the entire range of wind speeds from 00m/s to $30$m/s, weighted by the density of the estimated wind speeds from a. We use a wind-power curve of the form Carrillo et al. [48, Fig. 1]. This estimates the wind power that is generated for the respective wind speeds and accounts for the frequency at which these wind speeds occur. Finally, subplot c shows the cumulative sum of the generated wind powers over time. All wind-power values are normalized by the number of time steps on the respective temporal grids to align the wind-power scales. From the black lines aligning closely with the purple line, we conclude that the downscaled predictions preserve the aggregated wind-power generation as simulated by the ESM.
@@
Figure A.8: Local estimates of wind power over time. This plots investigates two pairs of locations that are each constraint by a single spatial node on the coarse climate grid. The layout of each row of the subplots (a) and (b) mirror figure A.7; each row corresponds to a single location on the fine-resolution grid, respectively.
A.5 Quantitative evaluation
We first apply quantile-mapping to bias-correct the climate patches. These frames are used as input for the different downscaling algorithms. Our model is compared to the following two baseline approaches.
Baseline 1: Conditioned frame-to-frame diffusion model
While our approach leverages a score-based approach, i.e., all states/trajectories are generated simultaneously in a non-autoregressive manner, we compare to a conventional frame-by-frame modeling approach using a denoising diffusion implicit model (DDIM) [5]. We use a time-conditional U-Net backbone, incorporating residual blocks and self-attention layers. The conditioning is then performed by bi-linearly upsampling the bias-corrected climate data to match the high-resolution data and concatenating them along the channel dimension. To predict the middle frame, we condition on a sequence of three frames [21].
Baseline 2: Bias-correction spatial disaggregation
The Bias Correction Spatial Disaggregation (BCSD) [49] serves as a baseline due to its simplicity and prominence in the downscaling community. The BCSD method consists of a quantile mapping to correct for biases, followed by bi-linear interpolation to match the spatial resolution of the reanalysis grid. To this end, reanalysis data is first downsampled and remapped to match the coarse resolution of the climate model grid. The bias correction step uses quantile mapping applied to the entire spatial domain. To account for seasonal variations, our implementation uses a moving average (with a window size of $25$ days) centered around each calender day when computing the quantiles. The temporal pooling ensures robust statistics by considering similar days from the seasonal cycle together.
@@ A.5 Quantitative evaluation
where $\Pi(P_{\text{pred}},P_{\text{ref}})$ is the set of couplings, that is, probability distributions whose marginals are $P_{\text{pred}}$ and $P_{\text{ref}}$. This variant makes the traditional Wasserstein distance computationally feasible for high-dimensional probability distributions [50]. In the sliced variant, the high-dimensional data is projected onto a one-dimensional line for which a one-dimensional Wasserstein distance can be efficiently computed. This slicing process is performed repeatedly for multiple different slices and the result is averaged to obtain a reliable metric.
The Mean Energy Log Ratio (MELR) is used to evaluate the preservation of potentially highly varying physical patterns in the downscaled spatial patches. The MELR is a metric that is derived from the radially averaged power spectral density (section A.2) as
@@
where $\mu_{x},\sigma_{x}$ and $\mu_{y},\sigma_{y}$ are the respective average and standard-deviations of the values in the respective patches $x$ and $y$. Analogously, $\sigma_{xy}$ is the covariance between the patches. Finally, $c_{1}$ and $c_{2}$ are constants to make the computation more robust. The SSIM takes on values between 00 and $1$, whereby higher values indicate more structural similarities througout the compared data points. An SSIM of $1$ can only be attained when both data points are identical. In table A.1 we report the SSIM using a window size of $k=15$.
Table A.1: Quantitative evaluation of downscaling methods. We compare the performance of our model, based on score-based data assimilation (SDA), to a conditional denoising diffusion implicit model (DDIM) and to bias-correction spatial disaggregation (BCSD). We report the sliced Wasserstein-1 distance (Sliced W1) over time, temporally averaged energy log ratio (MELR) and structural similarity index (SSIM) for the four variables mean sea-level pressure (psl), surface (2m) air temperature (tas), surface (10m) zonal uas and meridional (vas) wind speeds. The values are reported as mean $\pm$ standard deviation over the generated predictions. Note that, since the baseline methods are not capable of temporal downscaling, the quantitative evaluation is performed on the 6-hourly observation grid. Even though the method proposed in this work is under the additional constraint of generating temporally consistent 1-hourly trajectories, it beats both baselines in most cases. Between all approaches, the best performance per metric and variable is highlighted in bold.
@@
A.6 Approximating the gradient of the observation model
@@
$p(Y\mid X(\tau))$ $=\int p(Y\mid X(0),X(\tau))\leavevmode\nobreak\ p(X(0)\mid X(\tau))\leavevmode\nobreak\ \mathrm{d}X(0)$ (9)
$=\int p(Y\mid X(0))\leavevmode\nobreak\ p(X(0)\mid X(\tau))\leavevmode\nobreak\ \mathrm{d}X(0)$ (10)
@@
Linear stochastic differential equations, like the diffusion process defined in equation 1, can be equivalently formulated in terms of a discrete, linear Gaussian transition model
@@
for some time increment $\Delta\tau$. We refer to, e.g., Särkkä and Solin [34, Section 6.1] for more details on how to derive the transition and process-noise covariance functions $\mathrm{A}$ and $\mathrm{\Sigma}$ from the drift $\mathrm{F}$ and dispersion $\mathrm{L}$ of the diffusion process (equation 1). The forward diffusion process can be simulated over extended time ranges in a single step via sampling from the Gaussian transition density
@@
which is an equivalent formulation of equation 13 for the case $\tau=0$, i.e., starting from the noise-free data point [3, 4, 5]. The same is not possible in the reverse direction, which would make the generation process trivial. Unfortunately, the backwards transition model associated with equation 2 is Gaussian only for infinitesimally small time decrements $\Delta\tau\rightarrow 0$ [52, 31], which leads to sampling quality increasing with the number of steps used for simulating the generative process. However, it is possible to estimate the posterior mean $\hat{X}(0)$ based on the current score estimate $s_{\theta}(X(\tau),\tau)$. By re-arranging equation 14 we note that
@@
The first equality uses the direct correspondence between $\epsilon$ and the de-noising score function of equation 13 [3, 32]. Since the true score, like $\epsilon$, is not known while generating a new data point, it is in the next step approximated with our parametric score model $s_{\theta}$. A more detailed derivation can be found for the scalar case ($A,\Sigma\in\mathbb{R}$) in Chung et al. [42].
With that, we established the necessary background regarding the conditional model, in order to describe the approximation to the conditioning mechanism that is used in this work. Plugging into equations 7 and 8 the approximation from equations 9 to 12 reveals that we have to compute the gradient
@@
For our purposes, we assume a Gaussian observation model (equation 3), which simplifies this gradient to
@@
We plug the equality from equations 17 and 18 into the third component of the gradient to obtain
@@
The final step approximation avoids the computational demanding differentiation of the score function with respect to the perturbed state, which cuts away the majority of the computational cost in the conditioning. Altogether, the approximate conditioning term (c.f. equation 8), used throughout this work, is

### v4

@@ Methods
The code (in Python) for model training and evaluation, data processing, and for reproducing all presented experiments and figures contained in this manuscript, is available at https://github.com/schmidtjonathan/Climate2Weather. For the qunatile-mapping bias adjustment, the Python library python-cmethods [58] was used. Part of the plots were generated using the cartopy library [72]. The neural-network model and training are implemented using the PyTorch library [73].
Acknowledgements
The authors gratefully acknowledge financial support by the European Research Council through ERC CoG Action 101123955 ANUBIS; the DFG Cluster of Excellence "Machine Learning - New Perspectives for Science", EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); the DFG SPP 2298 (Project HE 7114/5-1), and the Carl Zeiss Foundation, (project "Certification and Foundations of Safe Machine Learning Systems in Healthcare"), as well as funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jonathan Schmidt and Luca Schmidt.
Author contribution statement
The joint project was initiated by J.S. and N.L. and coordinated by P.H. and N.L. J.S. implemented the code base for the model, training, experiments, data-processing and -pipeline (reanalysis data), as well as for evaluation and visualization of the results. J.S., L.S., and F.S. designed the experiments. L.S. implemented the data-processing for the climate simulations. The first version of the article was written by J.S. and L.S., after which all authors reviewed and edited the manuscript.
Competing Interests
The authors declare that they have no competing interests.
Supplementary Section 1 Spatial region: coarse and fine grid
Supplementary Figure 1: This plot shows the spatial region considered in this study. The coarse $8\times 8$-node grid is marked with purple crosses. The fine $128\times 128$-node grid is marked with green dots. Each coarse-grid node lies in the center of a corresponding $16\times 16$-patch of high-resolution grid nodes. Both grids span exactly the same area.
Supplementary Section 2 Spatial patterns on multiple scales
To assess the performance of the downscaling model across multiple spatial length scales, we show that the radially averaged power spectral densities (RAPSD) [74] of the predictions align with the reanalysis data in Supplementary Figure˜2. The RAPSD is computed by averaging the power spectrum over all directions of the same wavenumber in Fourier space. The quantity is commonly used in the context of weather dynamics, especially when estimating precipitation [48, 75, for example]. We use the open-source pysteps package by Pulkkinen et al. [76] to compute the RAPSD. The reported RAPSD values are averages over the considered time period.
@@ Acknowledgements
Supplementary Figure 2: This plot shows the RAPSD for reanalysis data, coarse input, and the corresponding downscaled predictions. Subplot (a) mirrors the experimental setup of the on-model experiment that predicts high-resolution reanalysis data during the cyclone "Friederike" in January 2018. Subplot (b) covers the CMIP6 downscaling setup as described in the Methods section.
@@
Supplementary Section 3 Embedding the predicted region in an extended spatial context
@@
We embed the high-resolution predictions during the cyclone "Friederike" (c.f. Figure 4) into a larger spatial context in order to learn about how long-distance interconnections between the studied spatial region and its surroundings are captured by the model. We argue that it is likely that our statistical downscaling model is able to capture the global connectedness of weather dynamics that are contained in the reanalysis data it is trained on. Concretely, Supplementary Figures 3 to 6 visualize the spatiotemporal dynamics of downscaled samples, reanalysis data, and conditioning input. Thereby, the spatial region is extended beyond the one considered in this study. Reanalysis data is used to fill in the regions outside of the predicted patch. Clearly visible or implausible transitions at the edges of the patch would indicate that the model predictions do not align with the surrounding spatial context. We find, however, that there is a smooth transition from outside the predicted area to its interior for the generated high-resolution predictions. The generated local dynamics seamlessly fit into the more surrounding context of reanalysis data.
@@
Supplementary Figure 3: Predicting mean sea-level pressure during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Supplementary Figure 4: Predicting surface air temperature during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Supplementary Figure 5: Predicting zonal wind speed during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
@@
Supplementary Figure 6: Predicting meridional wind speed during the cyclone "Friederike" (c.f. experiment above). Outside of the prediction range, which lies in the region indicated by red crosses, reanalysis data is filled in to give an impression as to how the predictions fit into more global dynamics.
Supplementary Section 4 Wind power prediction
@@
A crucial motivation for spatiotemporal downscaling climate simulations to the weather scale is downstream tasks that require future local weather patterns. We provide an exemplary evaluation of the estimated generated wind powers as computed from the CMIP6 simulation, the downscaled predictions, and the reanalysis data. We find that, when computing the spatial average of generated wind powers, the wind-speed and wind-power predictions of the ESM are matched by the downscaled samples in distribution (Supplementary Figure˜7). Comparing single locations reveals that the ESM locally sometimes over-predicts or under-predicts the wind power generated following the reanalysis data. For both cases, we consistently find that our model corrects the respective over- and under-estimation for multiple randomly selected locations. We compare estimated densities of the wind-speed values (Supplementary Figure˜7 a) and use this distribution to derive the amount of wind power generated from the respectively predicted wind speeds (Supplementary Figure˜7 b). For that, we first compute the wind-power curve for a range of wind-speed values (0m/s to 3030m/s), using the open-source package windpowerlib [77] and an arbitrary wind-turbine model (turbine type: "E-115/3000", hub-height: 100100m). Then, we weigh this power curve with the density of predicted wind speeds to obtain the predicted wind power for each wind speed, taking into account how likely this wind speed value is to occur in the predictions. Finally, we compare the accumulated generated wind power over time. We show the comparison between two pairs of locations in Supplementary Figure˜8.
@@
Supplementary Figure 7: Estimated generated wind power over time, averaged over the spatial domain. This plot analyzes the local wind-speed predictions via the generated wind-power derived from the wind speeds. The full spatial region is considered over the year 2014. Subplot a visualizes a kernel-density estimate of the aggregated wind speeds for reanalysis data (green), ESM simulations (purple), and downscaled predictions (black). In b, we plot the generated wind power for the entire range of wind speeds from 0m/s to 3030m/s, weighted by the density of the estimated wind speeds from a. We use a wind-power curve of the form Carrillo et al. [78, Fig. 1]. This estimates the wind power that is generated for the respective wind speeds and accounts for the frequency at which these wind speeds occur. Finally, subplot c shows the cumulative sum of the generated wind powers over time. All wind-power values are normalized by the number of time steps on the respective temporal grids to align the wind-power scales. From the black lines aligning closely with the purple line, we conclude that the downscaled predictions preserve the aggregated wind-power generation as simulated by the ESM.
@@
Supplementary Figure 8: Local estimates of wind power over time. This plot investigates two pairs of locations that are each constrained by a single spatial node on the coarse climate grid. The layout of each row of the subplots (a) and (b) mirrors Supplementary Figure˜7; each row corresponds to a single location on the fine-resolution grid, respectively.
Supplementary Section 5 Quantitative evaluation
We first apply quantile-mapping to bias-correct the climate patches. These frames are used as input for the different downscaling algorithms. Our model is compared to the following two benchmark approaches.
Benchmark 1: Interpolation
The first benchmark consists of a bias-correction step, followed by bi-linearly interpolating the coarse data to a finer spatial grid. This two-step procedure is called Bias Correction Spatial Disaggregation (BCSD) [79] and is both simple and prominently used in the downscaling literature. The reanalysis data is first remapped to match the coarse resolution of the climate model grid. The bias correction step then uses quantile mapping applied to the entire spatial domain. To account for seasonal variations, our implementation uses a moving average (with a window size of $25$ days) centered around each calendar day when computing the quantiles. The temporal pooling ensures robust statistics by considering similar days from the seasonal cycle together.
Benchmark 2: Conditioned frame-to-frame diffusion model
While our approach leverages a score-based approach, i.e., all states/trajectories are generated simultaneously in a non-autoregressive manner, we compare it to a conventional frame-by-frame modeling approach using a denoising diffusion implicit model (DDIM) [22]. We use a time-conditional U-Net backbone, incorporating residual blocks and self-attention layers. The conditioning is then performed by bi-linearly upsampling the bias-corrected climate data to match the high-resolution data and concatenating them along the channel dimension. To predict the middle frame, we condition on a sequence of three frames [45].
@@
where $\Pi(P_{\text{pred}},P_{\text{ref}})$ is the set of couplings, that is, probability distributions whose marginals are $P_{\text{pred}}$ and $P_{\text{ref}}$. This variant makes the traditional Wasserstein distance computationally feasible for high-dimensional probability distributions [80]. In the sliced variant, the high-dimensional data is projected onto a one-dimensional line for which a one-dimensional Wasserstein distance can be efficiently computed. This slicing process is performed repeatedly for multiple different slices, and the result is averaged to obtain a reliable metric.
The Mean Energy Log Ratio (MELR) is used to evaluate the preservation of potentially highly varying physical patterns in the downscaled spatial patches. The MELR is a metric that is derived from the radially averaged power spectral density (Section˜Supplementary Section 2) as
@@
where $\mu_{x},\sigma_{x}$ and $\mu_{y},\sigma_{y}$ are the respective average and standard-deviations of the values in the respective patches $x$ and $y$. Analogously, $\sigma_{xy}$ is the covariance between the patches. Finally, $c_{1}$ and $c_{2}$ are constants to make the computation more robust. The SSIM takes on values between 0 and $1$, whereby higher values indicate more structural similarities between the compared data points. An SSIM of $1$ can only be attained when both data points are identical. In Supplementary Table˜1 we report the SSIM using a window size of $k=15$.
Supplementary Table 1: Quantitative evaluation of downscaling methods. We compare the performance of our model, based on score-based data assimilation (SDA), to a conditional denoising diffusion implicit model (DDIM) and to bias-correction spatial disaggregation (BCSD). We report the sliced Wasserstein-1 distance (Sliced W1) over time, temporally averaged energy log ratio (MELR), and structural similarity index (SSIM) for the four variables: mean sea-level pressure (psl), surface (2m) air temperature (tas), surface (10m) zonal uas, and meridional (vas) wind speeds. The values are reported as mean $\pm$ standard deviation over the generated predictions. Note that, since the benchmark methods are not capable of temporal downscaling, the quantitative evaluation is performed on the 6-hourly observation grid. Even though the method proposed in this work is under the additional constraint of generating temporally consistent 1-hourly trajectories, it beats both benchmarks in most cases. Between all approaches, the best performance per metric and variable is highlighted in bold.
@@
Supplementary Section 6 Approximating the gradient of the observation model
@@
$p(Y\mid X(\tau))$ $=\int p(Y\mid X(0),X(\tau))\penalty 10000\ p(X(0)\mid X(\tau))\penalty 10000\ \mathrm{d}X(0)$ (11)
$=\int p(Y\mid X(0))\penalty 10000\ p(X(0)\mid X(\tau))\penalty 10000\ \mathrm{d}X(0)$ (12)
@@
Linear stochastic differential equations, like the diffusion process defined in Equation˜2, can be equivalently formulated in terms of a discrete, linear Gaussian transition model
@@
for some time increment $\Delta\tau$. We refer to, e.g., Särkkä and Solin [64, Section 6.1] for more details on how to derive the transition and process-noise covariance functions $\mathrm{A}$ and $\mathrm{\Sigma}$ from the drift $\mathrm{F}$ and dispersion $\mathrm{L}$ of the diffusion process (Equation˜2). The forward diffusion process can be simulated over extended time ranges in a single step via sampling from the Gaussian transition density
@@
which is an equivalent formulation of Equation˜15 for the case $\tau=0$, i.e., starting from the noise-free data point [20, 21, 22]. The same is not possible in the reverse direction, which would make the generation process trivial. Unfortunately, the backwards transition model associated with Equation˜3 is Gaussian only for infinitesimally small time decrements $\Delta\tau\rightarrow 0$ [82, 61], which leads to sampling quality increasing with the number of steps used for simulating the generative process. However, it is possible to estimate the posterior mean $\hat{X}(0)$ based on the current score estimate $s_{\theta}(X(\tau),\tau)$. By re-arranging Equation˜16 we note that
@@
The first equality uses the direct correspondence between $\epsilon$ and the de-noising score function of Equation˜15 [20, 62]. Since the true score, like $\epsilon$, is not known while generating a new data point, it is in the next step approximated with our parametric score model $s_{\theta}$. A more detailed derivation can be found for the scalar case ($A,\Sigma\in\mathbb{R}$) in Chung et al. [71].
With that, we established the necessary background regarding the conditional model, in order to describe the approximation to the conditioning mechanism that is used in this work. Plugging into Equations˜9 and 10 the approximation from Equations 11 to 14 reveals that we have to compute the gradient
@@
For our purposes, we assume a Gaussian observation model (Equation˜4), which simplifies this gradient to
@@
We plug the equality from Equations˜19 and 20 into the third component of the gradient to obtain
@@
The final step approximation avoids the computational demanding differentiation of the score function with respect to the perturbed state, which cuts away the majority of the computational cost in the conditioning. Altogether, the approximate conditioning term (c.f. Equation˜10), used throughout this work, is
@@
Supplementary Section 7 Anomalies
Supplementary Figure 9: Anomalies between downscaled and interpolated versions of coarse input. This plot supplements Figure 4. It visualizes the differences between a spatiotemporal interpolation of the coarse input $x_{\mathrm{interp}}$ (Figure 4, fourth row) and different fine-scale time series $\hat{x}$: 1) three samples from the proposed downscaling model (top three rows), 2) the average of the three samples (fourth row), and 3) the ground-truth reanalysis data (bottom row). This visualization exposes the local spatial and temporal patterns on the fine grid that are not contained in the coarse data. The rightmost column plots the corresponding temporal averages of the spatial anomalies. As in Figure 4, the downscaling model is only conditioned at 2PM (first column) and 8PM (penultimate column). Especially at those conditioning points, the local spatial patterns predicted in each sample (rows 1 through 3) are structurally similar to those in the reanalysis data (bottom row). Between the conditioning points (3PM through 7PM), the ground-truth anomalies (bottom row) expose that the temporal evolution of the cyclone is not predicted by the smooth temporal interpolation. Our model predicts spatiotemporal structure on the fine grid, adding information that cannot be trivially inferred from the coarse input.
Supplementary Section 8 Generative de-noising process
Supplementary Figure 10: The generative process of the diffusion model. Diffusion models learn a mapping from a tractable noise distribution ($\tau=1$; often Gaussian noise) to the training-data distribution. Using a statistical model for the score function, which separates noise from signal, an initial random-noise sample is iteratively de-noised into a data point that lies in a region of high data density. This figure shows seven (of 256) steps of this generative process, beginning at $\tau=1$ (top row; Gaussian noise) and ending at $\tau=0$ (bottom row). Fine-scale features are generated by the model towards the end of the generative process. Notably, the spatial and temporal structure emerge jointly, since the score function is estimated for the entire time series.
Supplementary Section 9 Relationship between downscaled variables
[…]
  1. 1.
    First, we draw from the prior. Sampling from the unconditioned generative downscaling model, yields an uninformed sequence of weather patterns.
  2. 2.
[…]
  3. 3.
    Third, we draw from the generative model that is conditioned only on v.
  4. 4.
[…]
Supplementary Figure˜11 visualizes these differently-informed samples of v as spatiotemporal sequences (cf. Figure 4). A comparison between the unconditioned sample (1.; first row in Supplementary Figure˜11), the different partly-conditioned predictions (2. and 3.; second and third rows in Supplementary Figure˜11), and the fully-conditioned predictions (4.; fourth row in Supplementary Figure˜11), demonstrates that the multivariate downscaling model has learned relationships between the variables, which it uses for generating downscaled predictions. Note that this experiment is purely diagnostic and serves to validate that the model learns and does not neglect relationships between the considered variables.
[…]